OpenStudy (anonymous):

In sunlight, a vertical stick 6 ft tall casts a shadow 2 ft long. At the same time a nearby tree casts a shadow 14 ft long. How tall is the tree? Round to the nearest tenth. 18.0 ft 85.7 ft 52.5 ft 42.0 ft

OpenStudy (jango_in_dtown):

 see tan theta=6/2=x/14 where x is the length of the tree then we have x=6*14/2=42 i.e. the tree is 42.0ft tall